Opened 10 years ago

Closed 8 years ago

Last modified 12 months ago

#233 closed flight processing (fixed)

GB08/02, flight day 078a/2009, Delamere Forest

Reported by: jatt Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/in_progress/2009/flight_data/uk/GB08_02-2009_078a_Delamere

Data arrived from ARSF via SATA N on 30th March 2009.

Scientific details: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-02 Application and Science Case.pdf

Priority: a4H

PI: Prof Mark Danson

Sensors:

  • Eagle (13/11/2009)
  • Hawk (13/11/2009)
  • ALS-II Lidar (22/03/2010)
  • Photography (07/01/2010)

Quicklook

TBD.

Change History (35)

comment:1 Changed 10 years ago by jatt

  • Owner mggr deleted

comment:2 Changed 10 years ago by mark1

Applanix has been processed and copied to asrf directory

comment:3 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:4 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:5 Changed 10 years ago by knpa

No boresight figures for this date (or indeed any prior to day 133 in '09), only given as 'variable'.

comment:6 Changed 10 years ago by knpa

  • Owner knpa deleted
  • Status changed from assigned to new

comment:7 Changed 10 years ago by knpa

  • Priority changed from alpha 5 to alpha 4 high

comment:8 Changed 10 years ago by mggr

  • Component changed from ARSF to Processing: general

comment:9 Changed 10 years ago by emca

  • Owner set to harg

comment:10 Changed 9 years ago by harg

Hawk flightline 4 (SWIR078-09-4) had no .hdr file. I have made a "fake". The line number is somewhat off: there should be a non-natural number of lines, so rounded down.

comment:11 Changed 9 years ago by harg

Lack of dark frames therefore radiometric values are an estimate rather than real data. Dark frames taken from hawk line 3.

comment:12 Changed 9 years ago by harg

The hyperspectral data is ready for delivery checking.

comment:13 Changed 9 years ago by mark1

  • Owner changed from harg to mark1

comment:14 Changed 9 years ago by mark1

Checked data located at ~airborne/workspace/GB08_02-2009_078a_Delamere/delivery/20091028/GB08_02/

Have removed txt and doc versions of logsheet and replaced with pdf in delivery.

Edited readme file to include project code, flight line numbers, typos corrected.

Example commands tested and work ok, all level 1 files open and look good.

Ready for delivery.

comment:15 Changed 9 years ago by harg

Boresight values for Eagle and Hawk:

      Pitch:       Roll:       Heading:
e1:   0.15         -0.63       0
e2:   0.07         -0.67       0.27
e3:   -0.1012467   -0.665408   -0.00117264571
e4:   -0.085899928 -0.605906   -0.000327635087

h1:   0.0639243    0.201953    -0.28800340
h2:   0.03092585   0.08278325  -0.92657365
h3:   -0.05        0.181       0
h4:   0.25         0.3         0

comment:16 Changed 9 years ago by harg

For Eagle and Hawk, GPT and SCT offsets both 0.0 in all cases.

comment:17 Changed 9 years ago by mark1

Data has been rsynced back to ~arsf/arsf_data/2009/flight_data/uk/GB08_02-2009_078a_Delamere/

comment:18 Changed 9 years ago by mark1

  • Owner changed from mark1 to harg

comment:19 Changed 9 years ago by harg

The Eagle and Hawk data for this project have been delivered this morning (13/11/09) to Professor Mark Danson of Salford University in three DVDs.

comment:20 Changed 9 years ago by mark1

  • Description modified (diff)

Digital Photography dispatched to M Danson on 07/01/2010

comment:21 Changed 9 years ago by mark1

rsynced photograph delivery to

~arsf/arsf_data/2009/flight_data/uk/GB08_02-2009_078a_Delamere/

comment:22 Changed 9 years ago by mark1

LiDAR data has been processed and is ready for delivery check at:
~airborne/workspace/GB08_02-2009_078a_Delamere/delivery/20100309/GB08_02

comment:23 Changed 9 years ago by emca

Delivery check complete. Loaded ascii files in LAG and the data looks good.
LIDAR data ready for delivery.

comment:24 Changed 9 years ago by mark1

  • Description modified (diff)

LIDAR data dispatched to PI on 22/3/2010

comment:25 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

LIDAR data rsynced to:
~arsf/arsf_data/2009/flight_data/uk/GB08_02-2009_078a_Delamere/

All data delivered, ready for archiving.

comment:26 Changed 9 years ago by knpa

  • Description modified (diff)
  • Owner changed from harg to knpa
  • Status changed from new to assigned

Started archiving.

Project missing event file (and rcd logfiles).

comment:27 Changed 9 years ago by knpa

Hyperspectral processing log files appear to be missing.

comment:28 Changed 9 years ago by knpa

Hawk raw line 4 log file appears to be missing.

comment:29 Changed 9 years ago by knpa

Prepared for archiving and tarball made available to NEODC.

comment:30 Changed 9 years ago by knpa

Downloaded and backed up by NEODC.

Moved to archive disk.

comment:31 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:32 Changed 6 years ago by tipo

Reprocessed nav and hyperspectral.

comment:33 Changed 2 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:34 Changed 21 months ago by dac

@asm re-uploaded reprocessed data to NEODC. Waiting for conformation data have been archived.

comment:35 Changed 12 months ago by asm

Archiving

All data archived at NEODC. Re-processed hyperspectral is slightly misplaced at :http://data.ceda.ac.uk/neodc/arsf/2009/GB08_02/GB08_02-2009_078a_Delamere/GB08_02-2009_078a_Delamere/GB08_02-078a-hyperspectral-20130724/ (a dir within a subdir with the same name). Will notify them but all data is there.

Note: See TracTickets for help on using tickets.